sketch the graphs of hte line and the parabola on a single set of axes. find the coordinates of any points of interstctions by solving algebraically.
2x+y=10, y=9-x^2

Answer by Alan3354(69443)

sketch the graphs of hte line and the parabola on a single set of axes. find the coordinates of any points of interstctions by solving algebraically.
------------
dl the FREE graph software at
http://www.padowan.dk
==================
2x+y=10, y=9-x^2
---
2x + (9-x^2) = 10
x^2 - 2x + 1 = 0
x = 1, x = 1
-----------
--> (1,8) the only intersection point.
The line is tangent to the parabola